OpenStudy (anonymous):

what's the general formula to find the vertex of an equation. for example: x=2y^2. I can draw the graph and see that the vertex is (0,0), but is there a mathematical way to figure it out?

OpenStudy (mathmate):

The general form of a quadratic equation is y=a(x-h)^2+k where (h,k) is the vertex. If the equation is not already in the above (canonical) form, you can convert to that form by completing the squares.

OpenStudy (anonymous):

Thank you

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

well actually this is a parabola that opens to the right, and cannot be written as \[y=a(x-k)^2+h\] but rather as \[x=a(y-k)^2+h\] but you are right, the vertex is (0,0)

OpenStudy (anonymous):

For a given quadratic y = ax2 + bx + c, the vertex (h, k) is found by computing h = –b/2a, and then evaluating y at h to find k. If you've already learned the Quadratic Formula, you may find it easy to memorize the formula for k, since it is related to both the formula for h and the discriminant in the Quadratic Formula: k = (4ac – b2) / 4a.

OpenStudy (anonymous):

at the risk of repeating myself this is \[x=2y^2\] and CANNOT

OpenStudy (anonymous):

Find the vertex of y = 3x2 + x – 2 and graph the parabola. To find the vertex, I look at the coefficients a, b, and c. The formula for the vertex gives me: h = –b/2a = –(1)/2(3) = –1/6 Then I can find k by evaluating y at h = –1/6: k = 3( –1/6 )2 + ( –1/6 ) – 2 = 3/36 – 1/6 – 2 = 1/12 – 2/12 – 24/12 = –25/12

OpenStudy (anonymous):

be written as \[y=a(x-k)^2+h\] because the square is on the "y" not on the "x"